# FIPS 204-Compatible Threshold ML-DSA via Masked Lagrange Reconstruction

### v4

@@ Abstract
We present the first threshold ML-DSA (FIPS 204) scheme achieving statistical share privacy (no computational assumptions) with arbitrary thresholds while producing standard 3.3 KB signatures verifiable by unmodified implementations. Our primary technique is Shamir nonce DKG: parties jointly generate the signing nonce via a distributed key generation protocol, so that both the nonce and the long-term secret are degree-(T1)(T\!-\!1) Shamir sharings. The honest party’s response has conditional min-entropy exceeding $5\times$ the secret key entropy, requiring no computational assumptions. In coordinator-based profiles (P1, P3+), this eliminates the two-honest requirement for response privacy ($|S|\geq T$ suffices); in the fully distributed profile (P2), the existing mask-hiding condition $|S\setminus C|\geq 2$ applies. As a secondary technique, we introduce pairwise-canceling masks ($|S\setminus C|\geq 2$) for commitment and r0-check values, addressing the three challenges unique to lattice-based threshold signing: the $\|\mathbf{z}\|_{\infty}$ rejection sampling check, secure evaluation of the r0-check without leaking $c\mathbf{s}_{2}$, and EUF-CMA security under the resulting Irwin-Hall nonce distribution.
We instantiate these techniques in three deployment profiles with complete UC proofs. Profile P1 uses a TEE coordinator (3 online rounds plus 1 offline preprocessing round, 6ms). Profile P2 eliminates hardware trust via MPC (5 rounds, 22ms for small thresholds). Profile P3+ uses lightweight 2PC with semi-asynchronous signing (2 logical rounds, 22ms), where signers precompute nonces offline and respond within a time window. Our Rust implementation scales from 2-of-3 to 32-of-45 thresholds with sub-100ms latency (P1 and P3+) and success rates of $\approx$21–45%, comparable to single-signer ML-DSA.
@@ 1.1 Motivation: The Threshold Gap
    Short-coefficient secret sharing [dPN25]: Uses carefully constructed linear secret sharing schemes (LSSS) where reconstruction coefficients remain small. This approach produces standard-size signatures and has been demonstrated for $T\leq 8$. The Ball-Çakan-Malkin bound [BCM21] shows that $\{0,1\}$-LSSS for threshold $T$ requires share size $\Omega(T\log T)$, which imposes practical limits on scaling this approach.
@@ 1.2 Our Contribution
We present Masked Threshold ML-DSA, the first threshold signature scheme for ML-DSA that achieves statistical share privacy (no computational assumptions) with arbitrary thresholds $T\leq N$ while producing standard-size signatures ($\approx 3.3$ KB for ML-DSA-65) identical to single-signer ML-DSA. Unlike prior pairwise-mask schemes that require a two-honest assumption ($|S|\geq T+1$), our protocol achieves response privacy with signing sets of size $|S|\geq T$ (the information-theoretic minimum for Shamir reconstruction) in coordinator-based profiles (P1, P3+), where individual masked commitments $\mathbf{W}_{i}$ are sent only to the coordinator rather than broadcast (Remark 3.7); commitment and r0-check mask hiding retains $|S\setminus C|\geq 2$ (Lemma 4.9), which applies to P2’s broadcast model. This addresses the threshold gap for $T\in[9,32]$ while remaining practical for configurations such as $(3,5)$, $(11,15)$, and $(16,21)$.
@@
The scheme achieves practical success rates comparable to single-signer ML-DSA (theoretical baseline $\approx 20\%$, observed $\approx$21–45% due to the more concentrated Irwin-Hall nonce distribution) and is fully verifier-compatible with FIPS 204. The aggregated nonce follows an Irwin-Hall distribution rather than uniform, but this does not affect EUF-CMA security and only introduces a quantifiable statistical difference analyzed in Section 3.
@@
Our approach is based on two techniques. The primary technique is Shamir nonce DKG: parties jointly generate the signing nonce via a distributed key generation protocol, producing a degree-(T1)(T\!-\!1) Shamir sharing of $\mathbf{y}$ that matches the structure of the long-term secret $\mathbf{s}_{1}$. Each party then computes $\mathbf{z}_{i}=\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i}$ without applying the Lagrange coefficient, and the combiner reconstructs $\mathbf{z}=\sum_{i}\lambda_{i}\mathbf{z}_{i}=\mathbf{y}+c\mathbf{s}_{1}$. Because the adversary observes only T1T-1 evaluations of the honest party’s degree-(T1)(T\!-\!1) polynomial, the remaining degree of freedom provides an approximate one-time pad: the honest party’s response $\mathbf{z}_{h}$ has conditional min-entropy exceeding $5\times$ the secret key entropy (for $|S|\leq 17$), requiring no computational assumptions.
@@
Our optimized Rust implementation achieves practical signing times: 5.8ms (P1-TEE), 21.5ms (P2-MPC), and 22.1ms (P3+-Async) for $(3,5)$-threshold on commodity hardware (including retries). P1 and P3+ remain under 100ms for all configurations up to $(32,45)$; P2 trades higher latency ($\sim$130–194ms for $T\geq 16$) for eliminating hardware trust. The communication cost is approximately 12.3 KB per party per signing attempt.
@@
We prove security in the random oracle model. Unforgeability reduces to the Module-SIS problem, matching the security of single-signer ML-DSA. Share privacy is statistical (no computational assumptions) via the Shamir nonce DKG, with conditional min-entropy exceeding $5\times$ the secret key entropy for $|S|\leq 17$; no computational assumptions are needed for response privacy. Commitment and r0-check privacy reduce to PRF security.
@@
A TEE/HSM coordinator performs the r0-check and hint generation inside a trusted enclave, achieving optimal 3-round signing. We prove EUF-CMA security under Module-SIS (Theorem 4.2), with statistical response privacy (Theorem 4.4) and PRF-based mask privacy (Lemma 4.9).
@@
A fully untrusted coordinator where the r0-check is computed via an MPC subprotocol using SPDZ [DPSZ12] and edaBits [EKMOZ20]. Using combiner-mediated commit-then-open and constant-depth comparison circuits, we achieve 5 online rounds (reduced from 8). We prove UC security against malicious adversaries corrupting up to N1N-1 parties (Theorem 6.2).
@@
Designed for human-in-the-loop authorization. Two designated Computation Parties (CPs) jointly evaluate the r0-check via 2PC, but signers participate semi-asynchronously: they precompute nonces offline and respond within a time window (e.g., 5 minutes) rather than synchronizing for multiple rounds. This achieves 2 logical rounds (1 signer step + 1 server round). We prove UC security under the assumption that at least one CP is honest (Theorem 6.5). The “Async” column in Table 1 distinguishes our work from concurrent approaches, which require synchronous multi-round signer participation.
@@ 1.4 Comparison with Prior and Concurrent Work
Bienstock et al. [BDEPT25] $\infty$ 3.3 KB 37–136 $\times$ $\times$ Comp.
@@
Trilithium [Trilithium25] 2 3.3 KB 3 $\times$ $\times$ Comp.
This work (P1) $\infty$ 3.3 KB 3 (TEE) $\times$ Stat.
This work (P2) $\infty$ 3.3 KB 5 $\times$ Stat.
This work (P3+) $\infty$ 3.3 KB 2 (1/2 CP) Stat.
Table 1: Comparison of threshold ML-DSA schemes. ✓= supported, $\times$ = not supported. “Dishon.” = dishonest-majority security (tolerates N1N-1 corruptions); Bienstock et al. require honest majority. “Privacy”: Comp. = computational (PRF-based), Stat. = statistical (no computational assumptions, min-entropy $\geq 5\times$ secret key entropy) for response shares $\mathbf{z}_{i}$; commitment/r0-check privacy uses PRF-based masks ($|S\setminus C|\geq 2$). “(TEE)” = dishonest-majority via TEE trust (not purely cryptographic); “(1/2 CP)” = via 1-of-2 computation-party honest assumption. P3+ = 1 signer step + 1 server round. P1 UC assumes ideal TEE (Theorem 4.13). Plus 1 offline preprocessing round (nonce DKG).
Three concurrent works address threshold ML-DSA with FIPS 204 compatibility. Bienstock et al. [BDEPT25] achieve arbitrary $T$ with UC security, but require honest-majority and 37–136 communication rounds. Celi et al. [CDENP26] achieve dishonest-majority with 6 rounds, but are limited to $T\leq 6$ due to the Ball-Çakan-Malkin bound [BCM21] on short-coefficient LSSS. Trilithium [Trilithium25] is UC-secure but restricted to 2 parties in a server-phone architecture.
@@
We contribute: (1) Shamir nonce DKG for threshold ML-DSA, achieving statistical response privacy (no computational assumptions) and in coordinator-based profiles (P1, P3+), eliminating the two-honest requirement for response shares ($|S|\geq T$ suffices; see Remark 3.7)—to our knowledge, the first threshold lattice-based signature with statistical share privacy and $|S|=T$ signing sets; (2) Pairwise-canceling masks adapted to the lattice setting for commitment and r0-check values, addressing rejection sampling and r0-check key leakage; (3) Three deployment profiles (P1-TEE, P2-MPC, P3+-2PC) with different trust/performance trade-offs, including the novel semi-async model for P3+; (4) Complete security proofs, with unforgeability reducing to Module-SIS, statistical privacy via Shamir nonce DKG (Theorem 4.4), and UC security proofs for all profiles; (5) A reference implementation with benchmarks from $(2,3)$ to $(32,45)$ threshold configurations.

### v5

# FIPS 204-Compatible Threshold ML-DSA via Shamir Nonce DKG

@@ Abstract
We present the first threshold ML-DSA (FIPS 204) scheme achieving statistical share privacy (no computational assumptions) with arbitrary thresholds while producing standard 3.3 KB signatures (ML-DSA-65, NIST Level 3) verifiable by unmodified implementations. Our primary technique is Shamir nonce DKG: parties jointly generate the signing nonce via a distributed key generation protocol, so that both the nonce and the long-term secret are degree-(T1)(T\!-\!1) Shamir sharings. This achieves statistical nonce share privacy (no computational assumptions): the honest party’s nonce share has conditional min-entropy exceeding $5\times$ the secret key entropy for $|S|\leq 17$ (tightest case: $5.02\times$ at $|S|=17$; statistical here means a min-entropy bound requiring no computational assumptions, not closeness to uniform in statistical distance—the bounded nonce range makes SD $\approx 0.9926$ from uniform inevitable). Key privacy of the aggregated signature—whether $\sigma$ reveals the signing key beyond the public key—follows under the same assumption as single-signer ML-DSA, which remains an open problem in the broader literature. In coordinator-based profiles (P1, P3+), this eliminates the two-honest requirement for nonce share privacy ($|S|\geq T$ suffices); in the fully distributed profile (P2), $|S\setminus C|\geq 2$ is required for commitment and r0-check mask-hiding (nonce share privacy itself holds at $|S|\geq T$ in P2 as well). As a secondary technique, we introduce pairwise-canceling masks ($|S\setminus C|\geq 2$) for commitment and r0-check values, addressing the three challenges unique to lattice-based threshold signing: the $\|\mathbf{z}\|_{\infty}$ rejection sampling check, secure evaluation of the r0-check without leaking $c\mathbf{s}_{2}$, and EUF-CMA security under the resulting Irwin-Hall nonce distribution.
We instantiate these techniques in three deployment profiles with complete UC proofs. Profile P1 uses a TEE coordinator (3 online rounds plus 1 offline preprocessing round, 5.8 ms for 3-of-5). Profile P2 eliminates hardware trust via MPC (5 rounds, 22ms for small thresholds). Profile P3+ uses lightweight 2PC with semi-asynchronous signing (2 logical rounds, 22ms), where signers precompute nonces offline and respond within a time window. Our Rust implementation scales from 2-of-3 to 32-of-45 thresholds with sub-100ms latency (P1 and P3+) and success rates of $\approx$21–45%, comparable to single-signer ML-DSA. A direct shift-invariance analysis (Corollary Q.24) shows the per-session EUF-CMA security loss from the Irwin-Hall nonce distribution is <0.007<0.007 bits for $|S|\leq 17$ and <0.013<0.013 bits for $|S|\leq 33$; over $q_{s}$ signing sessions the total loss is $0.013\cdot q_{s}$ bits (Corollary 4.3), eliminating the scalability gap found in prior work.
@@ 1.1 Motivation: The Threshold Gap
    Short-coefficient secret sharing [dPN25]: Uses carefully constructed linear secret sharing schemes (LSSS) where reconstruction coefficients remain small. This approach produces standard-size signatures and has been demonstrated for $T\leq 8$. Two separate arguments constrain scaling: (a) the Ball-Çakan-Malkin bound [BCM21] shows that $\{0,1\}$-LSSS for threshold $T$ requires share size $\Omega(T\log T)$, a communication complexity lower bound; (b) for ML-DSA specifically, the reconstruction coefficients of dPN25’s construction are binomial coefficients $\binom{T}{k}$, which exceed q=8,380,417q=8{,}380{,}417 when $T\geq 26$ (since $\binom{25}{12}=5{,}200{,}300<q$ but $\binom{26}{13}=10{,}400{,}600>q$; see Corollary Q.32), providing a tighter practical ceiling for this specific approach.
@@ 1.2 Our Contribution
We present Masked Threshold ML-DSA, the first threshold signature scheme for ML-DSA that achieves statistical share privacy (no computational assumptions) with arbitrary thresholds $T\leq N$ while producing standard-size signatures ($\approx 3.3$ KB for ML-DSA-65) syntactically identical in format to single-signer ML-DSA. Unlike prior pairwise-mask schemes that require a two-honest assumption ($|S|\geq T+1$), our protocol achieves nonce share privacy with signing sets of size $|S|\geq T$ (the minimum required by Shamir’s $(T,N)$ threshold property for reconstruction) in coordinator-based profiles (P1, P3+), where individual masked commitments $\mathbf{W}_{i}$ are sent only to the coordinator rather than broadcast (Remark 3.9); commitment and r0-check mask hiding retains $|S\setminus C|\geq 2$ (Lemma 4.15), which applies to P2’s broadcast model. This addresses the threshold gap for $T\in[9,32]$ while remaining practical for configurations such as $(3,5)$, $(11,15)$, and $(16,21)$.
@@
The scheme achieves practical success rates overlapping with single-signer ML-DSA (theoretical single-signer baseline $\approx 20$$25\%$; observed threshold range $\approx$21–45% (see Table 5); the higher end is due to the Irwin-Hall distribution’s greater concentration near zero, placing more nonces within the $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$ acceptance region relative to the uniform single-signer nonce) and is fully verifier-compatible with FIPS 204. The aggregated nonce follows an Irwin-Hall distribution rather than uniform, introducing a quantifiable security cost. We provide two analyses:
Conservative bound (Raccoon-style Rényi, Theorem 4.17): The per-session multiplicative advantage blowup is $\sqrt{R_{2}^{\mathsf{vec}}}\approx 2^{2.7}$ for $|S|\leq 17$, equivalent to an additive $\approx$2.7 bits of per-session security loss in bits (the advantage bound blows up by a factor $(R_{2}^{\mathsf{vec}})^{q_{s}/2}$, which corresponds to a linear $q_{s}\times 2.7$ bit loss when expressed in bits). This bound becomes vacuous for $q_{s}\geq 36$ queries and should not be used for concrete security estimates. At $|S|=33$, the per-session Rényi blowup satisfies $\log_{2}\sqrt{R_{2}^{\mathsf{vec}}}\approx 37.5$ bits; starting from the 96-bit post-Cauchy–Schwarz proof baseline (half the 192-bit ML-DSA-65 security parameter, arising from the Rényi-to-probability conversion), the provable security bound drops to ${\approx}58$ bits after a single signing session and is formally vacuous at $q_{s}=3$ ($96-3\times 38<0$). This is not a weakness of the scheme itself, but a fundamental limitation of the Raccoon-style bounding technique: the per-session Rényi divergence scales as $|S|^{4}$, making the proof unable to certify any security level for practical signing volumes at large $T$. This failure directly motivates the direct shift-invariance analysis introduced below.
[…]
@@
Our approach is based on two techniques. The primary technique is Shamir nonce DKG: parties jointly generate the signing nonce via a distributed key generation protocol, producing a degree-(T1)(T\!-\!1) Shamir sharing of $\mathbf{y}$ that matches the structure of the long-term secret $\mathbf{s}_{1}$. Each party then computes $\mathbf{z}_{i}=\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i}$ without applying the Lagrange coefficient, and the combiner reconstructs $\mathbf{z}=\sum_{i}\lambda_{i}\mathbf{z}_{i}=\mathbf{y}+c\mathbf{s}_{1}$. Because the adversary observes only T1T-1 evaluations of the honest party’s degree-(T1)(T\!-\!1) polynomial, the remaining degree of freedom provides an approximate one-time pad: the honest party’s nonce share $\mathbf{y}_{h}$ has conditional min-entropy exceeding $5\times$ the secret key entropy (for $|S|\leq 17$), requiring no computational assumptions (Theorem 4.7; proved).
Key privacy given the public signature is a distinct, open question: it is conjectured to hold computationally under the same Module-LWE assumption that secures ML-DSA itself, but a formal reduction from key privacy to Module-LWE remains unavailable for FIPS 204 ML-DSA specifically (with its HighBits/hint-vector structure), unresolved even in the single-signer FIPS 204 setting (open problem).
@@
Our Rust implementation achieves practical signing times: 5.8ms (P1-TEE), 21.5ms (P2-MPC), 22.1ms (P3+-Async) for (3,5)(3,\!5)-threshold on commodity hardware (per valid signature, 3344 expected attempts; $(2,3)$ achieves 4ms for P1 and 17ms for P3+, see Table 5). P1 and P3+ remain under 100ms per attempt for all configurations up to (32,45)(32,\!45); P2 trades higher latency (${\sim}130$194194ms per attempt for $T\geq 16$, non-monotone due to rejection-sampling variance) for eliminating hardware trust. Communication cost: $\approx$12.3 KB per party per attempt.
@@
We prove security in the random oracle model. Unforgeability reduces to the Module-SIS problem, matching the security of single-signer ML-DSA. Nonce share privacy is statistical (no computational assumptions) via the Shamir nonce DKG, with conditional min-entropy exceeding $5\times$ the secret key entropy for $|S|\leq 17$; key privacy given the public signature is conjectured to hold computationally under Module-LWE (open problem, unproven even for single-signer FIPS 204). Commitment and r0-check privacy reduce to PRF security.
@@
A TEE/HSM coordinator performs the r0-check and hint generation inside a trusted enclave, achieving optimal 3-round signing. We prove EUF-CMA security under Module-SIS (Theorem 4.2), with statistical nonce share privacy (Theorem 4.7) and PRF-based mask privacy (Lemma 4.15).
@@
A fully untrusted coordinator where the r0-check is computed via an MPC subprotocol using SPDZ [DPSZ12] and edaBits [EKMOZ20]. Using combiner-mediated commit-then-open and constant-depth comparison circuits, we achieve 5 online rounds (reduced from 8). We prove UC security against static malicious adversaries corrupting up to N1N-1 parties (Theorem 6.3).
@@
Designed for human-in-the-loop authorization. Two designated Computation Parties (CPs) jointly evaluate the r0-check via 2PC, but signers participate semi-asynchronously: they precompute nonces offline and respond within a time window (e.g., 5 minutes) rather than synchronizing for multiple rounds. This achieves 2 logical rounds (1 signer step + 1 server round). We prove UC security against static adversaries under the assumption that at least one CP is honest (Theorem 6.6). The “Async” column in Table 1 distinguishes our work from concurrent approaches, which require synchronous multi-round signer participation.
@@ 1.3 Technical Overview
Key algebraic observation.
The correctness of this construction rests on the fact that the partial response $\mathbf{z}_{i}=\mathbf{y}_{i}+c\mathbf{s}_{1,i}$ is $\mathbb{Z}_{q}$-linear in the Shamir evaluation. Because $\mathbf{y}$ and $\mathbf{s}_{1}$ are both degree-(T1)(T\!-\!1) Shamir polynomials over the same evaluation points, Lagrange interpolation distributes over addition:
$\sum_{i\in S}\lambda_{i}\mathbf{z}_{i}=\sum_{i\in S}\lambda_{i}\mathbf{y}_{i}+c\sum_{i\in S}\lambda_{i}\mathbf{s}_{1,i}=\mathbf{y}+c\mathbf{s}_{1}.$
This allows each signer to transmit the unscaled response $\mathbf{z}_{i}$—without pre-multiplying by the Lagrange coefficient $\lambda_{i}\in\mathbb{Z}_{q}$—so that $\|\mathbf{z}_{i}\|_{\infty}\leq\lceil\gamma_{1}/|S|\rceil+\beta$ holds at the individual level. The combiner applies $\{\lambda_{i}\}$ only at aggregation, where Lagrange interpolation recovers a valid $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$ satisfying $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$ via standard rejection sampling. By contrast, approaches based on generic LSSS with large reconstruction coefficients $\lambda_{i}\sim q$ must either pre-scale partial responses—producing individual norms $\|\lambda_{i}\mathbf{z}_{i}\|_{\infty}\sim q\gamma_{1}$, which universally violate the rejection bound and cause every signing attempt to abort—or employ noise flooding to mask the large shares, sacrificing FIPS 204 signature-size compatibility.
@@ 1.4 Comparison with Prior and Concurrent Work
Bienstock et al. [BdCE25] $\infty$ 3.3 KB 23–79a $\times$ $\times$ Comp.
@@
Trilithium [Trilithium25] 2 3.3 KB 14 $\times$ $\times$ Comp.
This work (P1) $\infty$ 3.3 KB $3{+}1^{\ddagger}$ (TEE) †§ $\times$ Stat.
This work (P2) $\infty$ 3.3 KB $5{+}1^{\ddagger}$ § $\times$ Stat.
This work (P3+) $\infty$ 3.3 KB 2 (1/2 CP) § Stat.
Table 1: Comparison of threshold ML-DSA schemes. ✓= supported, $\times$ = not supported. “Dishon.” = dishonest-majority unforgeability (tolerates N1N-1 corruptions for EUF-CMA; for P2, commitment/r0-check privacy additionally requires $|S\setminus C|\geq 2$, see ); Bienstock et al. require honest majority. “Privacy” column: Stat. = statistical nonce share privacy (no computational assumptions, conditional min-entropy $\geq 5\times$ secret key entropy) for nonce shares $\mathbf{y}_{i}$ conditioned on the DKG view only; key privacy given the public signature $\sigma$ is an open problem for this work (conjectured computational under M-LWE; unproven even for single-signer FIPS 204; see §4). Comp. for prior work = computational privacy guarantee per each scheme’s respective analysis (specifics vary: may cover nonce shares, response shares, or key leakage from the signature). Commitment/r0-check privacy for this work uses PRF-based masks ($|S\setminus C|\geq 2$ for P2). “(TEE)” = dishonest-majority via TEE trust (not purely cryptographic); “(1/2 CP)” = via 1-of-2 computation-party honest assumption. P3+ = 1 signer step + 1 server round. P1 UC assumes ideal TEE (Theorem 4.21)—a hardware trust assumption, not a cryptographic guarantee; P2’s dishonest-majority UC (Theorem 6.3) is purely cryptographic. Rounds shown as online++offline; the +1+1 offline nonce-DKG preprocessing round is structurally necessary to prevent rewinding and is amortizable over batches of signing sessions. P2 achieves dishonest-majority unforgeability (tolerates N1N-1 corruptions); commitment and r0-check privacy additionally requires $|S\setminus C|\geq 2$ (Lemma 4.15). At $T=N$ with N1N-1 corruptions, mask hiding does not hold in P2; P1 and P3+ retain nonce-share privacy at $|S\setminus C|=1$ via coordinator-only communication. §UC proofs for this work (Theorems 4.216.6) are against static adversaries (corruption set fixed before execution); adaptive adversary security is an open problem. aRound count for Bienstock et al. as reported in [BdCE25]; exact range depends on protocol variant and threshold $T$.
Three concurrent works address threshold ML-DSA with FIPS 204 compatibility. Bienstock et al. [BdCE25] achieve arbitrary $T$ with UC security, but require honest-majority and 23–79 communication rounds (range reflects threshold configurations from $T=2$ to $T=N$; per their protocol analysis). Borin et al. [BCDENP25] and Celi et al. [CDENP26] (the latter extending the former) achieve dishonest-majority with 6 rounds, but are limited to $T\leq 6$ by the constraints of their short-coefficient LSSS construction (the Ball-Çakan-Malkin bound [BCM21] establishes an asymptotic $\Omega(T\log T)$ lower bound on LSSS share size; their specific construction further requires small Lagrange reconstruction coefficients, limiting $T$ to single digits in practice). Trilithium [Trilithium25] is UC-secure but restricted to 2 parties in a server-phone architecture.
@@
We contribute: (1) Shamir nonce DKG for threshold ML-DSA, achieving statistical nonce share privacy (no computational assumptions) and in coordinator-based profiles (P1, P3+), eliminating the two-honest requirement for response shares ($|S|\geq T$ suffices; see Remark 3.9)—to our knowledge, the first threshold lattice-based signature achieving both: (i) statistical nonce share privacy (no computational assumptions), and (ii) $|S|=T$ signing sets in coordinator-based form. (Raccoon [Raccoon2024] achieves statistical nonce privacy via sum-of-uniform nonces without computational assumptions, but requires $|S|\geq T+1$ (not $|S|=T$) for nonce share privacy in its coordinator model, and targets Dilithium (not FIPS 204-compatible ML-DSA). Our improvement is specifically on the $|S|=T$ threshold; both schemes are non-computational for nonce privacy.); (2) Pairwise-canceling masks adapted to the lattice setting for commitment and r0-check values, addressing rejection sampling and r0-check key leakage; (3) Three deployment profiles (P1-TEE, P2-MPC, P3+-2PC) with different trust/performance trade-offs, including the novel semi-async model for P3+; (4) Complete security proofs, with unforgeability reducing to Module-SIS, statistical privacy via Shamir nonce DKG (Theorem 4.7), and UC security proofs for all profiles; (5) A reference implementation with benchmarks from $(2,3)$ to $(32,45)$ threshold configurations. Key generation currently assumes a trusted dealer for Shamir share distribution; a fully decentralized DKG achieving verifiable short-coefficient shares is an open problem discussed in Section 6.